OpenStudy (anonymous):

-16i in trigonometric form

OpenStudy (anonymous):

16(cos(π2)+isin(π2))

OpenStudy (anonymous):

-16i = 16 [cos 3pi/2 + i sin 3pi/2] So, by De Moivre's Theorem, the fourth roots are given by 16^(1/4) [cos((3pi/2 + 2 pi i k)/4) + i sin((3pi/2 + 2 pi i k)/4)], where k = 0,1,2,3. More specifically, they are k = 0 ==> 2 [cos(3pi/8) + i sin(3pi/8)] k = 1 ==> 2 [cos(7pi/8) + i sin(7pi/8)] k = 2 ==> 2 [cos(11pi/8) + i sin(11pi/8)] k = 3 ==> 2 [cos(15pi/8) + i sin(15pi/8)] --------------------------------------... 25i = 25 [cos pi/2 + i sin pi/2] So, by De Moivre's Theorem, the fourth roots are given by 25^(1/4) [cos((pi/2 + 2 pi i k)/4) + i sin((pi/2 + 2 pi i k)/4)], where k = 0,1,2,3. Note that 25^(1/4) = (5^2)^(1/4) = 5^(1/2). More specifically, the roots are k = 0 ==> sqrt(5) [cos(pi/8) + i sin(pi/8)] k = 1 ==> sqrt(5) [cos(5pi/8) + i sin(5pi/8)] k = 2 ==> sqrt(5) [cos(9pi/8) + i sin(9pi/8)] k = 3 ==> sqrt(5) [cos(13pi/8) + i sin(13pi/8)].

OpenStudy (anonymous):

Do u understand any of it?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

No probs

OpenStudy (anonymous):

Anymore?

OpenStudy (anonymous):

so I need to find the Square root of -16i…. I am using the theorem but i am just not coming up with the right answer.

OpenStudy (anonymous):

I dont know any other way to explain g2g be on later tho.

OpenStudy (anonymous):

4 -i

OpenStudy (anonymous):

So how can i put that in to trigonometric form?